# MMASD: A Multimodal Dataset for Autism Intervention Analysis

### v1

@@ Abstract.
Autism spectrum disorder (ASD) is a developmental disorder characterized by significant social communication impairments and difficulties perceiving and presenting communication cues. Machine learning techniques have been broadly adopted to facilitate autism studies and assessments. However, computational models are primarily concentrated on specific analysis and validated on private datasets in the autism community, which limits comparisons across models due to privacy-preserving data sharing complications. This work presents a novel privacy-preserving open-source dataset, MMASD as a MultiModal ASD benchmark dataset, collected from play therapy interventions of children with Autism. MMASD includes data from 32 children with ASD, and 1,315 data samples segmented from over 100 hours of intervention recordings. To promote public access, each data sample consists of four privacy-preserving modalities of data; some of which are derived from original videos: (1) optical flow, (2) 2D skeleton, (3) 3D skeleton, and (4) clinician ASD evaluation scores of children, e.g., ADOS scores. MMASD aims to assist researchers and therapists in understanding children’s cognitive status, monitoring their progress during therapy, and customizing the treatment plan accordingly. It also has inspiration for downstream tasks such as action quality assessment and interpersonal synchrony estimation. MMASD dataset can be easily accessed at https://github.com/Li-Jicheng/MMASD-A-Multimodal-Dataset-for-Autism-Intervention-Analysis.
multimodal dataset, machine learning, deep learning, autism spectrum disorder, human activity recognition
journalyear: 2023copyright: acmcopyrightconference: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION; October 9–13, 2023; Paris, Francebooktitle: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION (ICMI ’23), October 9–13, 2023, Paris, Franceprice: 15.00ccs: Social and professional topics People with disabilitiesccs: Human-centered computingccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Machine learning
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019) and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D/3D skeleton and clinician evaluation results. The dataset can be accessed publicly and investigate research questions centered on social, and behavioral interactions of children with ASD in playful, group activities.
@@
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019) and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D/3D skeleton and clinician evaluation results. The dataset can be accessed publicly and investigate research questions centered on social, and behavioral interactions of children with ASD in playful, group activities.
@@ 1. Introduction
Autism spectrum disorder (ASD) is a neurodevelopmental disorder characterized by significant impairments in social communication, as well as difficulties in perceiving and expressing communication cues. Approximately 1 in 54 children are on the spectrum in the United States, resulting in over 1 million affected individuals nationwide (Baio et al., 2018). Primary treatment of ASD includes behavioral and psycho-social interventions accompanied by prescribed medication. Behavioral and psycho-social interventions facilitate social and communication development, while medication helps control associated symptoms and comorbid problems (Billing et al., 2020). Specifically, psycho-social interventions are diverse in content and can vary in curriculum, structure, discipline, and theme. Prevailing therapeutic interventions include applied behavior analysis and robot-assisted therapy, both of which can provide valuable data for analyzing children’s mental development and developing individualized treatment plans.
Various studies have widely adopted machine learning techniques to facilitate autism research (Li et al., 2022a, b; Guo et al., 2023; Marinoiu et al., 2018; Pandey et al., 2020; Chen and Zhao, 2019; de Belen et al., 2020). Compared to traditional methods that rely heavily on human expertise and experience, machine learning approaches can help reduce the need for human labor and associated costs while still achieving decent performance. The autism community has benefited from machine learning techniques in many areas, including but not limited to autism diagnosis (Chen and Zhao, 2019; Wall et al., 2012), emotion recognition (Li et al., 2021b; Marinoiu et al., 2018), and movement pattern assessment (Li et al., 2022a, b, 2021a; Pandey et al., 2020).
@@
To overcome the data-sharing roadblock, in this paper, we proposed a publicly available multimodal dataset, MMASD [^1] MMASD can be accessed at https://github.com/Li-Jicheng/MMASD-A-Multimodal-Dataset-for-Autism-Intervention-Analysis., collected from physical therapy interventions of children with autism. MM-ASD maintains privacy while retaining essential movement features by providing optical flow, 2D and 3D skeletons that are derived from the original videos, thereby avoiding the exposure of sensitive and identifiable raw video footage. Additionally, it includes clinician evaluation results of each child such as ADOS-2 and motor function scores. We also provide the intervention activity class labels for overall scene understanding. Overall, MMASD can be used to help therapists and researchers to understand children’s cognitive status, track development progress in therapy, and guide the treatment plan accordingly. It also provides inspiration for downstream tasks, such as activity recognition (Pandey et al., 2020), action quality assessment (Li et al., 2021a), and interpersonal synchrony estimation (Li et al., 2022b). In contrast to current datasets, e.g., listed in Table  1, our dataset stands out for the following reasons:
@@
  • It includes multimodal features such as optical flow, 2D/3D skeletons, demographic data, and ADOS score rating. These features provide privacy-preserving approaches to maintain critical and full-body motion information.
@@ 20 ASD, 20  TD
1837 videos (83 frames/video)
@@ 32 ASD
Movement & Behavior analysis
Optical flow, 2D pose, 3D pose, ADOS score
@@ 2. Related Work
Another dataset related to humanoid robot interactions with autistic children is DE-ENIGMA (Riva et al., 2020), which includes using a multimodal human-robot interaction system to teach and expand social imagination in children with ASD. The DE-ENIGMA dataset comprises behavioral features such as facial mapping coordinates, visual and auditory, and facilitates communication and social interaction between the children and the robot. The authors indicated that the DE-ENIGMA could be used as an effective tool for teaching and expanding social imagination in children with ASD. They also suggest that the usage of a multimodal human-robot interaction could be a promising approach for developing interventions for children with ASD that aim to improve their social skills and promote better social integration.
@@
For action recognition dataset of children with ASD, Pandey et al. (Pandey et al., 2020) proposed a dataset of video recording actions and a technique to automate the response of video recording scenes for human action recognition. They evaluated their technique on two skill assessments with autism datasets and a real-world dataset of 37 children with ASD. Rehg et al. (Rehg et al., 2013) introduced a publicly available dataset including over 160 sessions of child-adult interactions. They discussed the use of computer vision and machine learning techniques to analyze and understand children’s social behavior in different contexts. They also identified technical challenges in analyzing various social behaviors, such as eye contact, smiling, and discrete behaviors. Rajagopalan et al. (Rajagopalan et al., 2013) explored the use of computer vision techniques to identify self-stimulatory behaviors in children with ASD. They also presented a self-stimulatory behavior dataset (SSBD) to assess the behaviors from video records of children with ASD in uncontrolled natural settings. The dataset comprises 75 videos grouped into three categories: arm flapping, head banging, and spinning behaviors.
@@ 3. Method
In the following sections, we describe the participants, procedure, and experimental settings of our proposed dataset.
@@ 3.1. Participants
We recruited 32 children (27 males and 5 females) with ASD from different races (Caucasian, African American, Asian, and Hispanic) and backgrounds through fliers posted online and onsite in local schools, services, and self-advocacy groups. Prior to enrollment, children were screened using the Social Communication Questionnaire (Srinivasan et al., 2016), and their eligibility was determined by the Autism Diagnostic Observation Schedule-2 (ADOS-2) (Lord et al., 2000; Srinivasan et al., 2015b) as well as clinical judgment. All the children were between 5 and 12 years old. Written parental consent was obtained from the Institutional Review Board of the university before enrollment. The Vineland Adaptive Behavior Scales (Sparrow and Cicchetti, 1989) were used to assess the children’s adaptive functioning levels. In general, 82% of the participating children had delays in the Adaptive Behavior Composite. Specifically, 70% of them experienced communication delays, 80% had difficulties with daily living skills, and 82% had delays in socialization
@@ 3.3. Experiment Settings
All videos were recorded in a house environment with the camera pointed toward the children. Models (trainers/therapists and trained adults) interacted with the child within a triadic context, and the adult model was the child’s confederate and practiced all activities with the child. Different tools were introduced to facilitate the training process depending on the theme of the intervention, for example, instruments and robots. Selected scenes in different themes of our proposed MMASD dataset are shown in Figure 2.
@@ 4. MMASD
  • Robot: children followed a robot and engaged in body movements.
@@
  • Yoga: children participated in yoga exercises led by experts. These exercises included body stretching, twisting, balancing, and other activities.
Overall, MMASD comprises 1,315 video clips that have been meticulously gathered from intervention video recordings spanning more than 108 hours. It consists of 244,679 frames with an average duration of 7.15 seconds. The average data length in MMASD is $7.0\pm 3.4$ seconds ($186.1\pm 92.9$ frames), with dimensions ranging from $320\times 240$ to $1920\times 1080$. Table 2 presents statistical information on MMASD. Depending on the activity conducted during the intervention, we further categorized all data into eleven activity classes as described in Table 3. Each activity class falls into a unique theme, as shown in Figure 2. MMASD also reports demographic and autism evaluation scores of all participating children, including date of birth, ADOS-2 score (social affect & restricted and repetitive behavior), motor functioning score, and severity of autism.
@@
The participant widens their knees as far as possible, places their feet with the big toes touching behind the buttocks, and assumes a shape like that of a frog in the kneeling position.
@@ 4.2. Data Annotation
In addition to the motion-related features mentioned above, we also reported clinician evaluation results such as ADOS-2 score, motor functioning score, and severity of autism for each participating child. ADOS-2 is a standardized assessment tool used to evaluate individuals suspected of having ASD. It is used in conjunction with other diagnostic information to help clinicians determine whether an individual meets the criteria for an ASD diagnosis. ADOS-2 includes several modules, each designed for individuals of different ages and language abilities and includes a series of activities and tasks that are used to observe key features of ASD, such as social communication skills and repetitive behaviors. The ADOS-2 scores are based on an individual’s performance during activities and tasks specific to their module and can range from 0 to 10 or higher depending on the algorithm used, with higher scores indicating more severe ASD symptoms. Moreover, we reported the ADOS comparison score, a continuous metric ranging from 1 to 10 that describes the severity of a child’s autism symptoms compared to children with ASD of similar age and language levels (Gotham et al., 2009). Low comparison scores are indicative of minimal evidence of autism symptoms, whereas high scores are indicative of severe autism symptoms.
@@ 4.4. Data Format
  • Demographic and clinical evaluation for ASD (9,)(9,)( 9 , ): including nine attributes such as participant ID, date of birth, chronological age, ADOS-2 module, social affect score, restricted and repetitive behavior score, ADOS comparison score, motor functioning score, and severity of autism, saved as CSV files.
@@ 4.5.1. Optical flow:
The Lucas-Kanade (Lucas and Kanade, 1981) method is used for our study. It is a popular technique used in computer vision to estimate the motion of objects between consecutive frames. The method assumes that the displacement of the image contents between two nearby instants is small and approximately constant within a neighborhood of the point under consideration. By solving the optical flow equation for all pixels within a window centered at the point, the method can estimate the motion of objects in the image sequence. Overall, the Lucas-Kanade optical flow method is an effective and popular technique for estimating motion in various computer vision applications.
@@
The OpenPose method (Cao et al., 2019) is used to extract 2D skeletons from our dataset of human action videos. OpenPose is a powerful tool for body, face, and hand analysis, developed by Carnegie Mellon University, and is based on deep learning techniques. It is a real-time multi-person key-point detection library that can accurately detect the key points of a human body, including joints and body parts, from an image or video feed. Initially, it predicts confidence maps for every body part and subsequently associates them with distinct individuals via Part Affinity Fields. The library is open-source and written in C++ with a Python API, which makes it easy to use and integrate into various computer vision applications.
@@
In our study, we employed the code and pre-trained model shared by the authors and used it on our dataset to suit our specific needs. By utilizing this method and applying it to our own data, we obtained 2D and 3D coordinates of key joints of the person(s).
@@ 5. Discussion
This section delves into the limitations of MMASD and the challenges we encountered during the dataset preparation process. As the experiments were conducted in real-world settings, we faced common computer vision challenges, including varying video quality, illumination changes, cluttered backgrounds, and pose variations. Notably, in the feature extraction stage, we encountered pose detection failures in challenging scenarios, such as body occlusion and participants moving out of the scene. The intrinsic video quality limitation of MMASD also restricted us from capturing subtle and fine-grained features, such as facial expressions.
Moreover, it is imperative to conduct in-depth investigations into domain-specific challenges. For instance, in standard benchmarks for human activity recognition, typically developing individuals exhibit dominant and continuous actions with similar intensity. However, data on MMASD may not solely contain the target behavior throughout the entire duration, as impromptu actions or distractions may occur during therapy sessions for children. Furthermore, unlike prevailing benchmarks that collect ground truth skeleton data by attaching sensors to the human body, MMASD generates skeleton data by means of pre-trained deep neural networks. This is because children with autism have limited tolerance for external stimuli, and the presence of sensors on their bodies may cause them to become anxious, agitated, or exhibit challenging behaviors. Consequently, the skeleton data’s reliability in MMASD depends on the performance of the underlying pose detectors.
In addition, children with autism can exhibit varying motor functions, resulting in different intensity levels and completion rates for the same activity. Clinician evaluation results can offer additional guidance in determining and classifying activities for each individual. Nevertheless, there is a need to explore ways to incorporate tabulated clinician evaluation results with other movement-related features to generate comprehensive representations.

### v3

@@ Abstract.
Autism spectrum disorder (ASD) is a developmental disorder characterized by significant impairments in social communication and difficulties perceiving and presenting communication signals. Machine learning techniques have been widely used to facilitate autism studies and assessments. However, computational models are primarily concentrated on very specific analysis and validated on private, non-public datasets in the autism community, which limits comparisons across models due to privacy-preserving data-sharing complications. This work presents a novel open source privacy-preserving dataset, MMASD as a MultiModal ASD benchmark dataset, collected from play therapy interventions for children with autism. The MMASD includes data from 32 children with ASD, and 1,315 data samples segmented from more than 100 hours of intervention recordings. To promote the privacy of children while offering public access, each sample consists of four privacy-preserving modalities, some of which are derived from original videos: (1) optical flow, (2) 2D skeleton, (3) 3D skeleton, and (4) clinician ASD evaluation scores of children. MMASD aims to assist researchers and therapists in understanding children’s cognitive status, monitoring their progress during therapy, and customizing the treatment plan accordingly. It also inspires downstream social tasks such as action quality assessment and interpersonal synchrony estimation. The dataset is publicly accessible via the MMASD project website.
multimodal dataset; machine learning; deep learning; autism spectrum disorder; human activity recognition; 2D/ 3D skeleton; privacy-preserving data sharing.
journalyear: 2023copyright: acmlicensedconference: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION; October 9–13, 2023; Paris, Francebooktitle: INTERNATIONAL CONFERENCE ON MULTIMODAL INTERACTION (ICMI ’23), October 9–13, 2023, Paris, Franceprice: 15.00doi: 10.1145/3577190.3614117isbn: 979-8-4007-0055-2/23/10submissionid: 1335ccs: Social and professional topics People with disabilitiesccs: Human-centered computingccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Machine learning
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019), and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D / 3D skeleton, and clinical evaluation results. The dataset is publicly accessible for addressing research questions centered on social and behavioral interactions of children with ASD in playful group activities.
@@
Figure 1. MMASD provides multiple multimodal privacy-preserving features derived from original videos via ROMP (Sun et al., 2021), OpenPose (Cao et al., 2019), and Lucas-Kanade (Lucas and Kanade, 1981), including optical flow, 2D / 3D skeleton, and clinical evaluation results. The dataset is publicly accessible for addressing research questions centered on social and behavioral interactions of children with ASD in playful group activities.
@@ 1. Introduction
Autism spectrum disorder (ASD) is a neurodevelopmental disorder characterized by significant impairments in social communication, as well as difficulties in perceiving and expressing communication cues. Approximately 1 in 54 children are on the spectrum in the United States, resulting in over 1 million affected individuals nationwide (Baio et al., 2018). Primary treatment of ASD includes behavioral and psychosocial interventions accompanied by prescribed medications. Behavioral and psychosocial interventions facilitate social and communication development, while medication helps control associated symptoms and comorbid problems (Billing et al., 2020). Specifically, psychosocial interventions are diverse in content and can vary in curriculum, structure, discipline, and theme. Prevailing therapeutic interventions include applied behavior analysis and robot-assisted therapy, both of which can provide valuable data for analyzing children’s mental development and developing individualized treatment plans.
Various studies have widely used machine learning techniques to facilitate autism research (Li et al., 2022a, b; Guo et al., 2023; Marinoiu et al., 2018; Pandey et al., 2020; Chen and Zhao, 2019; de Belen et al., 2020). Compared to traditional methods that rely heavily on human expertise and experience, machine learning approaches can help reduce the need for human labor and associated costs while still achieving decent performance. The autism community has benefited from machine learning techniques in many areas, including but not limited to autism diagnosis (Chen and Zhao, 2019; Wall et al., 2012), emotion recognition (Li et al., 2021b; Marinoiu et al., 2018), and movement pattern assessment (Li et al., 2022a, b, 2021a; Pandey et al., 2020).
@@
To overcome some of these ASD data-sharing challenges, we propose a publicly available multimodal ASD dataset, MMASD [^1] The MMASD Dataset is accessible via https://sites.udel.edu/hci-lab/mmasd-project/.. MMASD maintains privacy while retaining essential movement features by providing optical flow, 2D and 3D skeletons that are derived from the original play therapy videos, thereby avoiding the exposure of sensitive and identifiable raw video footage. Additionally, it includes clinician evaluation results of each child, such as motor function scores. We also provide the intervention activity class labels for overall scene understanding. Overall, MMASD can be used to help therapists and researchers to understand children’s cognitive status, track development progress in therapy, and guide the treatment plan accordingly. It also provides inspiration for downstream tasks, such as activity recognition (Pandey et al., 2020), action quality assessment (Li et al., 2021a), and interpersonal synchrony estimation (Li et al., 2022b). In contrast to current datasets, e.g., listed in Table  1, our dataset stands out for the following reasons:
@@
  • It includes multimodal features such as optical flow, 2D/3D skeletons, demographic, and clinical evaluation data. These features provide privacy-preserving approaches to maintain critical and full-body motion information.
@@ 20 ASD, 20  TD**
1,837 videos (83 frames/video)
@@ 32 ASD
Movement & social behavior analysis
Optical flow, 2D pose, 3D pose, Clinical score
@@ 2. Related Work
Another dataset related to humanoid robot interactions with ASD children is DE-ENIGMA (Riva et al., 2020), which includes using a multimodal human-robot interaction system to teach and expand social imagination among children with ASD. The DE-ENIGMA dataset comprises behavioral features such as facial mapping coordinates, visual and auditory, and facilitates communication and social interaction between the children and the robot. The authors indicated that the DE-ENIGMA could be used as an effective tool for teaching and expanding social imagination in children with ASD. They also suggest that the usage of a multimodal human-robot interaction could be a promising approach for developing interventions for children with ASD that aim to improve their social skills and promote better social integration.
@@
For the action recognition dataset of children with ASD, Pandey et al. (Pandey et al., 2020) proposed a dataset of video recording actions and a technique to automate the response of video recording scenes for human action recognition. They evaluated their technique on two skill assessments with autism datasets and a real-world dataset of 37 children with ASD. Rehg et al. (Rehg et al., 2013) introduced a publicly available dataset including over 160 sessions of child-adult interactions. They discussed the use of computer vision and machine learning techniques to analyze and understand children’s social behavior in different contexts. They also identified technical challenges in analyzing various social behaviors, such as eye contact, smiling, and discrete behaviors. Rajagopalan et al. (Rajagopalan et al., 2013) explored the use of computer vision techniques to identify self-stimulatory behaviors in children with ASD. They also presented a self-stimulatory behavior dataset (SSBD) to assess the behaviors from video records of children with ASD in uncontrolled natural settings. Their dataset comprised 75 videos grouped into three categories: arm flapping, head banging, and spinning behaviors.
@@ 3. Method
In the following sections, we describe the participants, procedure, and experimental settings of our proposed dataset. This study was approved by the University of Delaware’s Institutional Review Board (IRB) # $637082-12$.
@@ 3.1. Participants
We recruited $32$ children with ASD (27 males and 5 females) from different races (Caucasian, African American, Asian, and Hispanic) through flyers posted online and onsite in local schools, services, and self-advocacy groups. Prior to enrollment, children were screened using the Social Communication Questionnaire (Srinivasan et al., 2016), and their eligibility was determined by the Autism Diagnostic Observation Schedule-2 (ADOS-2) (Lord et al., 2000; Srinivasan et al., 2015b) as well as clinical judgment. All the children were between 5 and 12 years old. Written parental consent was obtained before enrollment. The Vineland Adaptive Behavior Scales (Sparrow and Cicchetti, 1989) were used to assess the children’s adaptive functioning levels. In general, 82% of the participating children had delays in the Adaptive Behavior Composite. Specifically, 70% of them experienced communication delays, 80% had difficulties with daily living skills, and 82% had delays in socialization
@@ 3.3. Experiment Settings
All videos from triadic settings were recorded in a house environment with the camera pointed toward the participating child. Different tools were introduced to facilitate the training process depending on the theme of the intervention, for example, instruments and robots. Selected scenes in different themes of our proposed MMASD dataset are shown in Figure 2.
@@ 4. MMASD Dataset
  • Robot: children followed a robot and imitated bodily movements.
@@
  • Yoga: children participated in yoga exercises led by therapists. These exercises included body stretching, twisting, balancing, and other activities.
Overall, MMASD comprises 1,315 video clips that have been meticulously gathered from intervention video recordings spanning more than 108 hours. It consists of 244,679 frames with an average duration of 7.15 seconds. The average data length in MMASD is $7.0\pm 3.4$ seconds ($186.1\pm 92.9$ frames), with dimensions ranging from $320\times 240$ to $1920\times 1080$. Table 2 presents statistical information on MMASD. Depending on the conducted activity during the intervention, we further categorized all data into eleven activity classes as described in Table 3. Each activity class falls into a unique theme, as shown in Figure 2. MMASD also reports demographic and autism evaluation scores of all participating children, including date of birth, motor functioning score, and severity of autism.
@@
The participant widens their knees as far as possible, places their feet with the big toes touching their body, and assumes a shape like that of a frog in the kneeling position.
@@ 4.2. Data Annotation
In addition to the motion-related features mentioned above, we also reported clinician evaluation results such as motor functioning score, and severity of autism for each participating child. ADOS-2 is a standardized assessment tool used to evaluate individuals suspected of having ASD. It is used in conjunction with other diagnostic information to help clinicians determine whether an individual meets the criteria for an ASD diagnosis. ADOS-2 includes several modules, each designed for individuals of different ages and language abilities and includes a series of activities and tasks that are used to observe key features of ASD, such as social communication skills and repetitive behaviors. The ADOS-2 scores are based on an individual’s performance during activities and tasks specific to their module and can range from 0 to 10 or higher depending on the algorithm used, with higher scores indicating more severe ASD symptoms. Moreover, we reported the ADOS comparison score, a continuous metric ranging from 1 to 10 that describes the severity of a child’s autism symptoms compared to children with ASD of similar age and language levels (Gotham et al., 2009). Low comparison scores are indicative of minimal evidence of autism symptoms, whereas high scores are indicative of severe autism symptoms.
@@ 4.4. Data Format
  • Demographic and clinical evaluation for ASD (9,)(9,)( 9 , ): including nine attributes such as participant ID, date of birth, chronological age, social affect score, restricted and repetitive behavior score, motor functioning score, and severity of autism, saved as CSV files.
@@ 4.5.1. Optical flow:
The Lucas-Kanade (Lucas and Kanade, 1981) method is used for our study. It is a popular technique used in computer vision to estimate the motion of objects between consecutive frames. The method assumes that the displacement of the image contents between two nearby instants is small and approximately constant within a neighborhood of the point under consideration. By solving the optical flow equation for all pixels within a window centered at the point, the method can estimate the motion of objects in the image sequence. Overall, the Lucas-Kanade optical flow method is an effective and preferred technique for estimating motion in various computer vision applications.
@@
The OpenPose method (Cao et al., 2019) is used to extract 2D skeletons from our dataset of human action videos. OpenPose is a powerful tool for body, face, and hand analysis, developed by Carnegie Mellon University, and is based on Convolutional Neural Networks (CNNs). It is a real-time multi-person key-point detection library that can accurately detect the key points of a human body, including joints and body parts, from an image or video feed. Initially, it predicts confidence maps for every body part and subsequently associates them with distinct individuals via Part Affinity Fields. The library is open-source and written in C++ with a Python API, which makes it easy to use and integrate into various computer vision applications.
@@
In our study, we employed the code and pre-trained model shared by the authors and used it on our dataset to suit our specific needs. By utilizing this method and applying it to our own data, we obtained 2D and 3D coordinates of key joints of the person(s). Since it is suitable for occluded scenes and noisy data, ROMP demonstrated its ability to successfully identify and represent the dynamics of our multi-class, multi-person activities.
@@ 5. Discussion
This section delves into the challenges, insights, and future opportunities of MMASD dataset. As the experiments were conducted in real-world settings in children’s homes, we faced common computer vision challenges, including varying video quality, illumination changes, cluttered backgrounds, and pose variations. Notably, in the feature extraction stage, we encountered pose detection failures in challenging scenarios, such as body occlusion and participants moving out of the scene. The intrinsic video quality limitation of MMASD also restricted us from capturing subtle and fine-grained features, such as facial expressions.
Moreover, it is imperative to conduct in-depth investigations into domain-specific challenges. For instance, in standard benchmarks for human activity recognition, typically developing individuals exhibit dominant and continuous actions with similar intensity. However, data on MMASD may not solely contain the target behavior throughout the entire duration, as impromptu actions or distractions may (and will) occur during therapy sessions for children with ASD. Furthermore, unlike prevailing benchmarks that collect ground truth skeleton data by attaching sensors to the human body, MMASD generates skeleton data by means of pre-trained deep neural networks. This is because children with autism have limited tolerance for external stimuli, and the presence of sensors on their bodies may cause them to become anxious, agitated, or exhibit challenging behaviors. Consequently, the skeleton data’s reliability in MMASD depends on the performance of the underlying pose detectors. In addition, children with autism can exhibit varying motor functions, resulting in different intensity levels and completion rates for the same activity.
There are several directions for future work that are worth exploring. Firstly, further research can be conducted to develop and compare machine learning models on the MMASD dataset for various tasks, such as action quality assessment (Li et al., 2021a), interpersonal synchrony estimation (Li et al., 2022a, b), and cognitive status tracking (Li et al., 2023; Chheang et al., 2023). This can help establish benchmark performance and identify state-of-the-art methods for analyzing the full-body movements of children during therapeutic interventions. In addition, new approaches can be investigated to overcome pose detection failures in MMASD. For example, by introducing pose uncertainty (Li et al., 2022b) or attention mechanism to assign higher weights to more reliable body joints. Furthermore, the MMASD dataset can be expanded in several aspects. This includes new features such as mutual gaze (Guo et al., 2023), or additional annotations including movement synchrony scores or task-specific clinician evaluations. Also, the MMASD dataset can be expanded by the inclusion of (age- and gender-matched) typically developing children data for developing more comprehensive and contrastive models. Finally, efforts can be made to dataset augmentation via existing benchmarks not limited to the autism domain by matching samples with similar motion features (Pandey et al., 2020), which can significantly expand the scale of the autism dataset.